ROTATIONAL SPECTRA OF THE Ne-NH_{3} and Kr-NH_{3} VAN DER WAALS COMPLEXES

\underline{JENNIFER VAN WIJNGAARDEN} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

Rotational spectra of the Ne-NH_{3} and Kr-NH_{3} van der Waals complexes were measured between 4 and 24 GHz using a pulsed jet cavity Fourier transform microwave spectrometer. The isotopomers studied include those of ^{14}NH_{3} and ^{15}NH_{3} with ^{22}Ne and ^{20}Ne and with the 5 most abundant isotopes of Kr. Transitions corresponding to the RG+NH_{3} Σ(j=0, |k| = 0) internal rotor state have been measured and assigned for the Ne and Kr isotopomers and transitions related to the Σ(j=1, |k| = 1) state have been tentatively assigned for the Kr isotopomers. The spectra of the complexes containing 83Kr and ^{14}N showed nuclear hyperfine structure due to the quadrupole moments of the nuclei. Rotational, centrifugal distortion and quadrupole coupling constants were determined and used to calculate structural parameters.